a man is grilling on his 35.1 meter high balcony when the wind blows an aluminum serving tray at 22.33-degree Celcius out of his hands, sending it crashing to the ground. if all of the heat generated by this process is transferred to the tray. what will the tray's final temperature be? (the specific capacity of aluminum is 890 J/kg-K)
Answer #1

given that

height h = 35.1 m

initial temperature T1 = 22.33 C = 22.33+273.15 = 295.48 K

final temperature T2 = ?

specific heat capacity of aluminium c = 890 J/kg k

here gravitational potential energy is converted to the heat energy of aluminium tray

heat energy = mc(T2-T1)

gravitational potential energy = mgh

mc(T2-T1) = mgh

T2-T1 = gh/c

T2-295.48 = 9.8*35.1/890 = 0.386

T2 = 295.866 K = 22.716 C